# Deep Insights into Cognitive Decline: A Survey of Leveraging Non-Intrusive Modalities with Deep Learning Techniques

### v1

@@ Abstract
Cognitive decline is a natural part of aging, often resulting in reduced cognitive abilities. In some cases, however, this decline is more pronounced, typically due to disorders such as Alzheimer’s disease. Early detection of anomalous cognitive decline is crucial, as it can facilitate timely professional intervention. While medical data can help in this detection, it often involves invasive procedures. An alternative approach is to employ non-intrusive techniques such as speech or handwriting analysis, which do not necessarily affect daily activities. This survey reviews the most relevant methodologies that use deep learning techniques to automate the cognitive decline estimation task, including audio, text, and visual processing. We discuss the key features and advantages of each modality and methodology, including state-of-the-art approaches like Transformer architecture and foundation models. In addition, we present works that integrate different modalities to develop multimodal models. We also highlight the most significant datasets and the quantitative results from studies using these resources. From this review, several conclusions emerge. In most cases, the textual modality achieves the best results and is the most relevant for detecting cognitive decline. Moreover, combining various approaches from individual modalities into a multimodal model consistently enhances performance across nearly all scenarios.
Index Terms:
Cognitive Decline, Deep Learning, Data Modality, Non-Intrusive Techniques
@@ I Introduction
Cognitive abilities play a crucial role in daily life, constituting a fundamental aspect of our interactions with others. These abilities include reasoning, various memory functions, attention, and language comprehension and communication [1]. As individuals age, cognitive function typically undergoes a natural decline, with certain abilities, such as processing speed, reasoning, and memory, showing more pronounced signs of deterioration, while others remain relatively unaffected. [2]. Despite cognitive decline is a common and natural aspect of aging, some individuals may experience a more accelerated or pronounced deterioration in these abilities due to cognitive disorders, such as dementia [3]. The severity of this deterioration spans a broad spectrum, with Mild Cognitive Impairment (MCI) representing an intermediate stage between typical age-related decline and more severe forms, including different types of dementia [4, 5]. The decline in cognitive abilities due to aging represents a significant challenge for our society, particularly in light of the trend toward an aging population [6]. Nonetheless, not all disorders affecting cognitive abilities are associated with age. One such disorder is aphasia, which is usually caused by strokes [7].
@@
Recent technological advances, particularly deep learning, have the potential to revolutionize numerous fields, including healthcare. The integration of automation into medical practice offers promising opportunities to increase the efficiency and productivity of healthcare professionals [8, 9]. In the context of cognitive decline estimation, these technologies have demonstrated potential in facilitating early detection [10, 11], allowing timely intervention in the early stages of the disorder. Early detection is crucial, as it enables patients to receive professional treatment from the onset of the disorder. This timely intervention is particularly important, as previous studies have shown that support from doctors, psychiatrists, or psychologists can significantly improve patients’ quality of life and help mitigate cognitive decline over time, ultimately benefiting their cognitive and behavioral functioning [12].
@@
Deep learning methods are known to be data-hungry and can benefit from multiple data sources, each of which presents its own trade-off between contribution to detection, cost, privacy, and degree of intrusiveness to the patient. Many studies rely on medical data, including Magnetic Resonance Images (MRI) of patients’ brains, to diagnose these conditions [13, 14, 15, 16, 17, 18, 19]. Despite providing relevant information on the disorders and their effects on patient’s brains, medical data collection is notably invasive. Additionally, these approaches face limitations in terms of feasibility in different contexts. As a result, this study focuses on non-invasive techniques that avoid disrupting patients’ daily routines. These alternative approaches rely on video recordings, integrating visual, audio, and textual data. In this context, patient privacy, ease of deployment, and ensuring minimal interference with daily activities are key factors.
@@
Although various reviews and surveys have already addressed cognitive decline estimation, most rely solely on medical imaging, such as MRI or EHR (Electronic Health Records) [20, 21, 22]. Imaging methods have been also studied specifically for certain conditions, including Alzheimer’s or MCI [23, 24, 25, 26, 27]. Similarly, Alzheimer’s or Parkinson’s have also been investigated using non-intrusive modalities [28, 29, 30]. However, we find a notable deficit in the literature regarding comprehensive studies on overall cognitive decline estimation that leverage non-intrusive data modalities. This work aims to fill this gap in the literature.
@@
In this study, we propose a survey focused on the early prediction of cognitive decline using deep learning methods and non-intrusive data. To achieve this, we systematically review existing methodologies in this field and introduce the most relevant datasets for training and evaluating the performance of models designed to predict cognitive decline. Additionally, we discuss the effectiveness of each modality and identify the most suitable approaches for utilizing them. Furthermore, we explore various strategies for combining these modalities into multimodal models to leverage their unique capabilities and features. Employing these modalities enables a more comprehensive understanding of the issue, ultimately leading to improved performance on a variety of tasks. In summary, our main contributions are as follows:
@@
  • To the best of our knowledge, this is the first survey focused on the general estimation of cognitive decline using non-intrusive modalities with deep learning methods. The modalities included in this survey are audio, text, video, and image.
@@
  • We discuss the effectiveness and characteristics of each methodology, both for individual modalities and multimodal fusions. This analysis allows us to identify the most promising challenges and future research directions. Furthermore, we highlight the key issues and difficulties related to cognitive decline estimation tasks.
@@ III Review Methodology
By applying these quality assessment questions, we can ensure that the final works included in our study are of high quality and suitable for further analysis.
@@ IV Datasets
TABLE I: Cognitive disorders estimation datasets. For simplicity, “AD” denotes Alzheimer’s disease, “PAD” denotes probable Alzheimer’s disease, and “DTP” denotes different types of dementia.
@@
In this study, we have included basic metrics such as accuracy and F1-Score for the classification of cognitive impairment in the aforementioned datasets. Additionally, we have reported the primary metric used for a regression task. This task involves predicting a subject’s cognitive score based on the MMSE, and is measured by the Root Mean Square Error (RMSE). These metrics are applied in the selected works to ensure consistency in performance evaluation.
@@ V-A Audio Modality
The audio modality represents how humans perceive and process auditory information from the environment, and is typically captured by microphones. This modality is easy to obtain and, in the case of recording speech, can provide valuable insights into how we speak, as well as facilitate the extraction of other crucial modalities, such as text by a transcription process. Besides text transcription, there is also a wide range of features that can be representative, particularly in how we express our thoughts. The way we express our ideas can be very significant in cognitive estimation, as people with more severe decline tend to struggle to find the correct and appropriate words, leading to long pauses, hesitation, and the use of incorrect words among other issues [68, 69]. Furthermore, speech is frequently employed to identify an individual’s emotional state, a process known as Speech Emotion Recognition (SER) [70]. SER typically uses a range of speech features, including prosody, pitch, and rhythm to accurately predict emotions such as happiness or fear [71, 72]. Some studies have shown that cognitive disorders negatively affect emotional mood [73, 74], making it another relevant source of information for this task.
@@
DementiaBank Pitt Corpus [90] 2021 Audio Features 82.59 82.94 -
@@
Alzheimer & MCI NCMMSC2021 L [78] 2023 Transformers 85.70 85.70 -
@@
Aphasia Mandarin AphasiaBank [98] 2022 2D Spectrograms - - 3.53
V-A1 2D Spectorgrams Methods
@@
Another notable work using Mel Spectrograms is presented in [91, 128] and [92], both focusing on the classification of Alzheimer’s using the DementiaBank Pitt Corpus. In the case of [91], Mel Spectrograms were combined with pre-trained CNNs DenseNet [129], MobileNet [130], and ResNet [131], with the best results achieved using pre-trained DenseNet (accuracy of 73.49%). In the case of [92], the proposed pre-trained CNNs included ResNet18 [131], ResNet34 [131], ResNet50 [131], SqueezeNet [132], and VGG16 [133], with the best results obtained using VGG16, achieving an accuracy of 66% and an F1-Score of 62%. When combined with demographic data from patients, the accuracy and F1-Score both improved to 73%.
In addition to the use of CNNs to extract features from images, there are also studies where the embeddings obtained from these CNNs are fed into a LSTM model. The LSTM model is a specific type of recurrent neural network that enhances the capabilities of previous recurrent models by integrating long-term dependencies for sequential data [134]. The work of [75] employs this method to predict Alzheimer’s using the ADReSSo dataset, achieving an accuracy of 78.9%. In this study, the employed model is the pre-trained VGG16 model. The same method has been used in the work of [98] to predict the severity of Aphasia using the Mandarin AphasiaBank dataset. In this work the pre-trained ResNet model has been used as the CNN. This application to Aphasia achieved an RMSE score of 3.53 on the cognitive score of the subjects. This study also proposes the accuracy in determining each level of severity of Aphasia, but not for the differentiation between aphasic and normal cognitive subjects.
@@ V-A3 Audio Features Methods
There are different works, such as those proposed by L. Zhaoci et al. [90] and S. Zhengyan et al. [96], where both use pre-trained models for Automatic Speech Recognition (ASR) and their intermediate hidden states in combination with CNNs and LSTMs to predict Alzheimer’s. In [90], this methodology achieved an accuracy of 82.59% and an F1-Score of 82.94% on the DementiaBank Pitt Corpus dataset. In [96], the best result was an accuracy of 82.18% on their own dataset, predicting both Alzheimer’s and MCI.
@@ V-A4 Transformer Methods
As previously stated, Transformer models are employed in different modalities, including audio. This architecture typically comprises two principal components: an encoder and a decoder. The encoder processes the given input, resulting in embeddings, which serve as a representation of data. After being processed, embeddings capture the most relevant information. With this representation, the decoder part of the Transformer generates new data based on these representations [167, 168]. In the problem presented in this study, we are addressing a classification task for identifying a disorder and a regression task for estimating cognitive state. Therefore, we will focus exclusively on the encoder component, which is responsible for obtaining the data representation and not for generating new data. This representation allows for the prediction of either a disorder or the cognitive state.
@@
Regarding the cognitive estimation task, several works utilize the Transformer architecture to process audio signals. One of the most notable is proposed by Z. Cui et al. [76], which not only predicts Alzheimer’s disease using the ADReSSo dataset but also examines the impact of depression on this disease. Some studies link emotional mood, particularly depression, with cognitive impairment disorders. A significant percentage of people suffering from Alzheimer’s exhibit neuropsychiatric symptoms, and half of those with Alzheimer’s also suffer from depression [174]. Additionally, studies show that individuals with a history of depression have a higher risk of developing Alzheimer’s [175, 176]. Therefore, this study also incorporates a depression dataset, specifically the DAIC-WOZ dataset [177]. For this study, various Transformer-based architectures were employed, including WavLM [178], HuBert [179], and Wav2Vec2 [118]. Among these, the WavLM model performed best, achieving an F1-Score of up to 80.5, slightly surpassing the others. In addition to these proposed models, a fine-tuned version of WavLM using the depression dataset was also tested, but this approach did not improve upon the basic WavLM model. Another proposed method involved using a hidden state from the transcription model Whisper [123], but this did not yield improvements.
@@
Another study that employed Wav2Vec2 is the one proposed by Y. Ying et al. [78], which works with three different datasets for Alzheimer’s and MCI recognition. In this study, the Transformer-based model outperforms the extraction of various audio features, including eGeMAPS and IS10_lingusitc. This model achieves an accuracy of up to 71.2% and an F1-Score of 73.1% on the ADReSSo dataset for Alzheimer’s classification. It also attains an accuracy and F1-Score of 85.7% on the NCMMSC2021 long speech dataset and 81.7% and 81%, respectively, on the NCMMSC2021 short speech dataset, both for the classification of Alzheimer’s, MCI, and healthy subjects.
Similarly, the work in [95] utilizes the Wav2Vec2 model in combination with a dense layer to predict MCI on the I-CONECT dataset, achieving up to 50.42% accuracy and 54.02% F1-Score. The poor performance of this methodology is because the most relevant features in this dataset come from the video modality.
@@ V-B Text Modality
Recent works have achieved results similar to those of professional transcribers [123]. However, when applied to the domain of health, there are additional challenges that must be addressed. For instance, in the context of cognitive decline, uncertainty or indecision may impede the accurate transcription of audio. Nevertheless, there are some works that address this issue [181, 182, 183]. In this context, the work [184] uses the aforementioned Transformer-based Wav2Vec model to accurately transcribe the corresponding audio, incorporating pauses for further prediction of dementia using textual information.
@@
DementiaBank Pitt Corpus [91] 2023 Transformer 90.36 - -
@@
Alzheimer & MCI NCMMSC2021 L [78] 2023 Transformer 70.70 70.00 -

### v2

@@ Abstract
Cognitive decline is a natural part of aging. However, under some circumstances, this decline is more pronounced than expected, typically due to disorders such as Alzheimer’s disease. Early detection of an anomalous decline is crucial, as it can facilitate timely professional intervention. While medical data can help, it often involves invasive procedures. An alternative approach is to employ non-intrusive techniques such as speech or handwriting analysis, which do not disturb daily activities. This survey reviews the most relevant non-intrusive methodologies that use deep learning techniques to automate the cognitive decline detection task, including audio, text, and visual processing. We discuss the key features and advantages of each modality and methodology, including state-of-the-art approaches like Transformer architecture and foundation models. In addition, we present studies that integrate different modalities to develop multimodal models. We also highlight the most significant datasets and the quantitative results from studies using these resources. From this review, several conclusions emerge. In most cases, text-based approaches consistently outperform other modalities. Furthermore, combining various approaches from individual modalities into a multimodal model consistently enhances performance across nearly all scenarios.
@@ I Introduction
Cognitive abilities play a crucial role in daily life and are fundamental to our interactions with others. These abilities include reasoning, memory functions, attention, and language comprehension and communication [1]. As individuals age, cognitive function typically undergoes a natural decline, with some abilities, such as processing speed, reasoning, and memory, showing more pronounced deterioration, while others remain relatively unaffected [2]. Although cognitive decline is a common and natural aspect of aging, some individuals may experience a more accelerated or pronounced deterioration in these abilities due to cognitive disorders, such as dementia [3]. The severity of this deterioration spans a broad spectrum, with Mild Cognitive Impairment (MCI) representing an intermediate stage between typical age-related decline and more severe forms, including different types of dementia [4, 5]. Age-related cognitive decline presents a significant societal challenge, especially given the growing trend toward aging populations [6]. Nonetheless, not all disorders affecting cognitive abilities are associated with age. For instance, aphasia, typically caused by strokes, can impair language functions across age groups [7].
@@
Recent technological advances, particularly in deep learning, have the potential to revolutionize numerous fields, including healthcare. The integration of automation into medical practice offers promising opportunities to improve the efficiency and productivity of healthcare professionals [8, 9]. In the context of cognitive decline detection, these technologies have demonstrated potential in facilitating early detection [10, 11], enabling timely intervention during the initial stages. Early detection is crucial, as it allows patients to receive professional treatment from the onset. This timely intervention is particularly important, as previous studies have shown that support from doctors, psychiatrists, or psychologists can significantly improve patients’ quality of life and slow the progression of cognitive decline, ultimately benefiting their cognitive and behavioral functioning [12].
@@
These deep learning methods typically require large amounts of data and can benefit from multiple data sources, each offering a different balance between diagnostic contribution, cost, privacy, and degree of intrusiveness to the patient. Many studies rely on medical data, including Magnetic Resonance Images (MRI) of patients’ brains, to diagnose these conditions [13, 14, 15, 16, 17, 18, 19]. While this type of data offers valuable insights into disorders and their impact on patients’ brains, it typically requires highly invasive collection methods. Furthermore, these approaches face limitations in terms of feasibility in different contexts. As a result, this study focuses on non-invasive techniques that avoid disrupting patients’ daily routines. These alternative approaches often involve the use of video recordings, integrating visual, audio, and textual data. In this context, key priorities include protecting patient privacy, ensuring ease of deployment, and minimizing interference with everyday activities.
@@
Although various reviews and surveys have already addressed cognitive decline detection, most focus primarily on medical imaging, such as MRI or EHR (Electronic Health Records) [20, 21, 22]. Imaging methods have been also explored specifically for certain conditions, including Alzheimer’s or MCI [23, 24, 25, 19, 26]. In parallel, non-intrusive modalities have been used to study Alzheimer’s and Parkinson’s [27, 28, 29]. However, there remains a notable gap in the literature regarding comprehensive studies focused on general cognitive decline detection using non-intrusive data sources. This work aims to address that gap.
@@
In this study, we present a comprehensive survey focused on the early detection of cognitive decline using deep learning methods and non-intrusive data sources. To achieve this, we systematically review current methodologies in this field and introduce the most relevant datasets for training and evaluating the performance of models designed to detect cognitive decline. Additionally, we discuss the effectiveness of each modality and identify the most suitable approaches for leveraging them. Furthermore, we explore various strategies for combining these modalities into multimodal models to exploit their complementary characteristics and features. Employing these modalities enables a more comprehensive understanding of the issue, ultimately leading to improved performance on a variety of tasks. This survey is primarily intended for researchers in machine learning, particularly those developing deep learning models for healthcare applications. However, we also aim to present clinically relevant insights to support interdisciplinary collaboration with healthcare professionals interested in cognitive assessment.
In summary, our main contributions are as follows:
@@
  • To the best of our knowledge, this is the first survey focused on the general detection of cognitive decline using non-intrusive modalities with deep learning methods. The modalities included in this survey are audio, text, video, and image.
@@
  • We discuss the effectiveness and characteristics of each methodology, both for individual modalities and multimodal fusions. This analysis allows us to identify the most promising challenges and future research directions. Furthermore, we highlight the key issues and difficulties related to cognitive decline detection tasks.
@@ III Review Methodology
By applying these quality assessment questions, we can ensure that the final works included in our study are of high quality and suitable for further analysis. Each study has been evaluated using these questions, with responses scored as follows: Yes (1 point), Partially (0.5 points), and No (0 points). After scoring all candidate studies, an inclusion cutoff threshold of 3.5 out of 6 points was established. Based on this criterion, 43 studies were deemed of sufficient quality and included in the final analysis.
@@ IV Datasets
TABLE I: Cognitive disorders estimation datasets. For simplicity, “V” denotes Video modality, “A” denotes Audio modality, “T” denotes Text modality, “AD” denotes Alzheimer’s disease, “PAD” denotes probable Alzheimer’s disease, and “DTP” denotes different types of dementia.
@@
In this study, we have included basic metrics such as accuracy and F1-Score for the detection of cognitive impairment in the aforementioned datasets. Additionally, we have reported the primary metric used for a regression task. This task involves predicting a subject’s cognitive score based on the MMSE, and is measured by the Root Mean Square Error (RMSE). These metrics are applied in the selected works to ensure consistency in performance evaluation.
@@ V-A Audio Modality
The audio modality represents how humans perceive and process auditory information from the environment, and is typically captured by microphones. This modality is easy to obtain and, in the case of recording speech, can provide valuable insights into how we speak, as well as facilitate the extraction of other crucial modalities, such as text by a transcription process. Besides text transcription, there is also a wide range of features that can be representative, particularly in how we express our thoughts. The way we express our ideas can be very significant in cognitive estimation, as people with more severe decline tend to struggle to find the correct and appropriate words, leading to long pauses, hesitation, and the use of incorrect words among other issues [67, 68]. Furthermore, speech is frequently employed to identify an individual’s emotional state, a process known as Speech Emotion Recognition (SER) [69]. SER typically uses a range of speech features, including prosody, pitch, and rhythm to accurately detect emotions such as happiness or fear [70, 71]. Some studies have shown that cognitive disorders negatively affect emotional mood [72, 73], making it another relevant source of information for this task.
@@
DementiaBank
Pitt Corpus
[89] 2021 Audio Features 82.59 82.94 -
@@
Alzheimer
& MCI
NCMMSC2021 L [77] 2023 Transformers 85.70 85.70 -
@@
Aphasia
Mandarin
AphasiaBank
[97] 2022 2D Spectrograms - - 3.53
V-A1 2D Spectrograms Methods
@@
Another notable work using Mel Spectrograms is presented in [90, 127] and [91], both focusing on the detection of Alzheimer’s using the DementiaBank Pitt Corpus. In the case of [90], Mel Spectrograms were combined with pre-trained CNNs DenseNet [128], MobileNet [129], and ResNet [130], with the best results achieved using pre-trained DenseNet (accuracy of 73.49%). In the case of [91], the proposed pre-trained CNNs included ResNet18 [130], ResNet34 [130], ResNet50 [130], SqueezeNet [131], and VGG16 [132], with the best results obtained using VGG16, achieving an accuracy of 66% and an F1-Score of 62%. When combined with demographic data from patients, the accuracy and F1-Score both improved to 73%.
In addition to the use of CNNs to extract features from images, there are also studies where the embeddings obtained from these CNNs are fed into a LSTM model. The LSTM model is a specific type of recurrent neural network that enhances the capabilities of previous recurrent models by integrating long-term dependencies for sequential data [133]. The work of [74] employs this method to detect Alzheimer’s using the ADReSSo dataset, achieving an accuracy of 78.9%. In this study, the employed model is the pre-trained VGG16 model. The same method has been used in the work of [97] to detect the severity of Aphasia using the Mandarin AphasiaBank dataset. In this work the pre-trained ResNet model has been used as the CNN. This application to Aphasia achieved an RMSE score of 3.53 on the cognitive score of the subjects. This study also proposes the accuracy in determining each level of severity of Aphasia, but not for the differentiation between aphasic and normal cognitive subjects.
@@ V-A3 Audio Features Methods
There are different works, such as those proposed by L. Zhaoci et al. [89] and S. Zhengyan et al. [95], where both use pre-trained models for Automatic Speech Recognition (ASR) and their intermediate hidden states in combination with CNNs and LSTMs to detect Alzheimer’s. In [89], this methodology achieved an accuracy of 82.59% and an F1-Score of 82.94% on the DementiaBank Pitt Corpus dataset. In [95], the best result was an accuracy of 82.18% on their own dataset, detecting both Alzheimer’s and MCI.
@@ V-A4 Transformer Methods
As previously stated, Transformer models are employed in different modalities, including audio. This architecture typically comprises two principal components: an encoder and a decoder. The encoder processes the given input, resulting in embeddings, which serve as a representation of data. After being processed, embeddings capture the most relevant information. With this representation, the decoder part of the Transformer generates new data based on these representations [166, 167]. In the problem presented in this study, we are addressing a classification task for identifying a disorder and a regression task for estimating cognitive state. Therefore, we will focus exclusively on the encoder component, which is responsible for obtaining the data representation and not for generating new data. This representation allows for the detection of either a disorder or the cognitive state.
@@
Regarding the cognitive estimation task, several works utilize the Transformer architecture to process audio signals. One of the most notable is proposed by Z. Cui et al. [75], which not only detects Alzheimer’s disease using the ADReSSo dataset but also examines the impact of depression on this disease. Some studies link emotional mood, particularly depression, with cognitive impairment disorders. A significant percentage of people suffering from Alzheimer’s exhibit neuropsychiatric symptoms, and half of those with Alzheimer’s also suffer from depression [173]. Additionally, studies show that individuals with a history of depression have a higher risk of developing Alzheimer’s [174, 175]. Therefore, this study also incorporates a depression dataset, specifically the DAIC-WOZ dataset [176]. For this study, various Transformer-based architectures were employed, including WavLM [177], HuBert [178], and Wav2Vec2 [117]. Among these, the WavLM model performed best, achieving an F1-Score of up to 80.5, slightly surpassing the others. In addition to these proposed models, a fine-tuned version of WavLM using the depression dataset was also tested, but this approach did not improve upon the basic WavLM model. Another proposed method involved using a hidden state from the transcription model Whisper [122], but this did not yield improvements.
@@
Another study that employed Wav2Vec2 is the one proposed by Y. Ying et al. [77], which works with three different datasets for Alzheimer’s and MCI recognition. In this study, the Transformer-based model outperforms the extraction of various audio features, including eGeMAPS and IS10_lingusitc. This model achieves an accuracy of up to 71.2% and an F1-Score of 73.1% on the ADReSSo dataset for Alzheimer’s detection. It also attains an accuracy and F1-Score of 85.7% on the NCMMSC2021 long speech dataset and 81.7% and 81%, respectively, on the NCMMSC2021 short speech dataset, both for the detection of Alzheimer’s, MCI, and healthy subjects.
Similarly, the work in [94] utilizes the Wav2Vec2 model in combination with a dense layer to detect MCI on the I-CONECT dataset, achieving up to 50.42% accuracy and 54.02% F1-Score. The poor performance of this methodology is because the most relevant features in this dataset come from the video modality.
@@ V-B Text Modality
Recent works have achieved results similar to those of professional transcribers [122]. However, when applied to the domain of health, there are additional challenges that must be addressed. For instance, in the context of cognitive decline, uncertainty or indecision may impede the accurate transcription of audio. Nevertheless, there are some works that address this issue [180, 181, 182]. In this context, the work [183] uses the aforementioned Transformer-based Wav2Vec model to accurately transcribe the corresponding audio, incorporating pauses for further detection of dementia using textual information.
@@
DementiaBank
Pitt Corpus
[90] 2023 Transformer 90.36 - -
@@
Alzheimer
& MCI
NCMMSC2021 L [77] 2023 Transformer 70.70 70.00 -